Show that the following equation is exact
( x² + y²) dx + 2xydy = 0
1
Expert's answer
2020-03-18T15:51:00-0400

We known that, The necessary and sufficient condition for the differential equation

"Mdx+Ndy=0"

where "M" and "N" are function of "x" and "y", to be exact is that

"\frac {\partial M} {\partial y} =\frac {\partial N}{\partial x}" .

According to the question,

"M=x^2+y^2" and "N=2xy"

Therefore,"\frac {\partial M}{\partial y}=2y" and "\frac {\partial N}{\partial x}=2y" .

Hence, the given equation "(x^2+y^2)dx+2xydy=0"

is exact.